# Separable commutative algebras in equivariant homotopy theory

### v1

@@ 10. Not all separable algebras are standard
Consider $\mathbb{S}_{G}\in\operatorname{Sp}_{G}$, and note that $\pi_{0}^{G}(\mathbb{S}_{G})=A(G)$ the Burnside ring of the finite group $G$. By a result of Dress [Dre69], the ring $A(G)$ is indecomposable if and only if $G$ is solvable. It follows that if $G$ is not solvable then $\mathbb{S}_{G}$ decomposes non-trivially as a direct product $\mathbb{S}_{G}[f^{-1}]\times\mathbb{S}_{G}[(1-f)^{-1}]$ of separable algebras, which we claim are not standard. Indeed suppose that $\mathbb{S}_{G}[f^{-1}]\simeq\mathbb{D}(X_{+})$ for some finite $G$-set $X$. Since $\Phi^{e}\mathbb{S}_{G}=\mathbb{S}\in\operatorname{Sp}$ is indecomposable, we must have either $\Phi^{e}\mathbb{S}_{G}[f^{-1}]=0$ or $\Phi^{e}\mathbb{S}_{G}[(1-f)^{-1}]=0$. In the first case we deduce that $\Phi^{e}\mathbb{D}(\Sigma_{+}^{\infty}X)=0$ so $X=\varnothing$. This is a contradiction as we assumed that $\mathbb{S}_{G}[f^{-1}]\not=0$. In the second case we must have $\Phi^{e}\mathbb{S}_{G}[f^{-1}]=\mathbb{S}_{G}$. By checking on $e$-geometric fixed points we deduce that $X=\{\ast\}$, so that $\mathbb{S}_{G}[f^{-1}]\simeq\mathbb{S}_{G}$. This is a contradiction as $1-f$ is a nontrivial idempotent on $\mathbb{S}_{G}$ but trivial on $\mathbb{S}_{G}[f^{-1}]$.
@@
Pick $[b^{\prime},c^{\prime}]\in\pi_{0}(\mathscr{B})\times_{\pi_{0}(\mathscr{D})}\pi_{0}(\mathscr{C})$ so that $[f(b^{\prime})]=[g(c^{\prime})]$, that is, there exists an isomorphism $\eta\colon f(b^{\prime})\xrightarrow{\sim}g(c^{\prime})$ in $\mathscr{D}$. By definition of pullback in groupoids, the triple $(b^{\prime},c^{\prime},\eta^{\prime})$ defines an object $a^{\prime}\in\mathscr{A}$ which satisfies $\varphi(a^{\prime})=[b^{\prime},c^{\prime}]$, showing surjectivity of $\varphi$. For the second claim consider $a^{\prime}=[b^{\prime},c^{\prime},\eta^{\prime}\colon f(b^{\prime})\xrightarrow{\sim}g(c^{\prime})]\in\pi_{0}(\mathcal{A})$ and suppose that $\varphi(a)=\varphi(a^{\prime})$ which means that there exist isomorphisms $\beta\colon b\xrightarrow{\sim}b^{\prime}$ and $\gamma\colon c\xrightarrow{\sim}c^{\prime}$. We define a map (of sets)
@@
[…]
@@
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(\mathbb{1}^{X},\mathbb{1}^{X})$ $=\coprod_{X}\prod_{X}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(\mathbb{1},\mathbb{1})$
$=\coprod_{X}\prod_{X}\operatorname{Map}_{\operatorname{Fin}}(\ast,\ast)$
@@
Let $G=C_{6}$ be the cyclic group of order $6$. We take the following exhaustive collection of families
$\varnothing\subseteq\{e\}\subseteq\{e,C_{2}\}\subseteq\{e,C_{2},C_{3}\}\subseteq\mathrm{All}.$
@@
${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Fun}(BG,\operatorname{Sp}^{\omega}))}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{2}^{\wedge}G))\times\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{3}^{\wedge}G)).}$$\scriptstyle{L^{\omega}_{\{e\}}}$$\scriptstyle{(\widetilde{\Phi}^{e})^{\omega}}$$\scriptstyle{\beta\circ(\overline{\Phi}^{e})^{\omega}}$$\scriptstyle{\beta\circ\pi}$
Using 9.6 and the fact that the corresponding Weyl groups are either $2$-groups or $3$-groups, we see that
$\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})\simeq(\mathcal{F}\mathrm{in}_{G}/\{e\})^{\operatorname{op}},$
i.e., all separable algebras in $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})$ are standard. Using Lemma 7.6, we also find that
@@
We also observe that $(\mathbb{S}_{2}^{\wedge})^{tG}$ and $(\mathbb{S}_{3}^{\wedge})^{tG}$ are indecomposable (argue as in the proof of Theorem 5.5), so the stable module categories appearing in the bottom right corner have indecomposable unit. By Proposition 10.2 any object $A\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})$ can be represented by a tuple
@@
    an object $A_{1}\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})$;
@@
    an isomorphism $\eta_{1}\colon A_{1}^{\prime}\simeq A_{2}^{\prime}$ between the images of $A_{1}$ and $A_{2}$ in $\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{2}^{\wedge}G))$;
@@
    an isomorphism $\eta_{2}\colon A_{1}^{\prime\prime}\simeq A_{2}^{\prime\prime}$ between the images of $A_{1}$ and $A_{2}$ in $\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{3}^{\wedge}G))$.

### v2

@@ 10. Not all separable algebras are standard
Consider $\mathbb{S}_{G}\in\operatorname{Sp}_{G}$, and note that $\pi_{0}^{G}(\mathbb{S}_{G})=A(G)$ the Burnside ring of the finite group $G$. By a result of Dress [DRE69], the ring $A(G)$ is indecomposable if and only if $G$ is solvable. It follows that if $G$ is not solvable then $\mathbb{S}_{G}$ decomposes non-trivially as a direct product $\mathbb{S}_{G}[f^{-1}]\times\mathbb{S}_{G}[(1-f)^{-1}]$ of separable algebras, which we claim are not standard. Indeed suppose that $\mathbb{S}_{G}[f^{-1}]\simeq\mathbb{D}(X_{+})$ for some finite $G$-set $X$. Since $\Phi^{e}\mathbb{S}_{G}=\mathbb{S}\in\operatorname{Sp}$ is indecomposable, we must have either $\Phi^{e}\mathbb{S}_{G}[f^{-1}]=0$ or $\Phi^{e}\mathbb{S}_{G}[(1-f)^{-1}]=0$. In the first case we deduce that $\Phi^{e}\mathbb{D}(\Sigma_{+}^{\infty}X)=0$ so $X=\varnothing$. This is a contradiction as we assumed that $\mathbb{S}_{G}[f^{-1}]\not=0$. In the second case we must have $\Phi^{e}\mathbb{S}_{G}[f^{-1}]=\mathbb{S}_{G}$. By checking on $e$-geometric fixed points we deduce that $X=\{\ast\}$, so that $\mathbb{S}_{G}[f^{-1}]\simeq\mathbb{S}_{G}$. This is a contradiction as 1f1-f is a nontrivial idempotent on $\mathbb{S}_{G}$ but trivial on $\mathbb{S}_{G}[f^{-1}]$.
@@
Pick $b^{\prime},c^{\prime}]\in\pi_{0}(\mathscr{B})\times_{\pi_{0}(\mathscr{D})}\pi_{0}(\mathscr{C})$ so that $f(b^{\prime})]=[g(c^{\prime})]$, that is, there exists an isomorphism $\eta\colon f(b^{\prime})\xrightarrow{\sim}g(c^{\prime})$ in $\mathscr{D}$. By definition of pullback in groupoids, the triple $(b^{\prime},c^{\prime},\eta^{\prime})$ defines an object $a^{\prime}\in\mathscr{A}$ which satisfies $\varphi(a^{\prime})=[b^{\prime},c^{\prime}]$, showing surjectivity of $\varphi$. For the second claim consider $a^{\prime}=[b^{\prime},c^{\prime},\eta^{\prime}\colon f(b^{\prime})\xrightarrow{\sim}g(c^{\prime})]\in\pi_{0}(\mathcal{A})$ and suppose that $\varphi(a)=\varphi(a^{\prime})$ which means that there exist isomorphisms $\beta\colon b\xrightarrow{\sim}b^{\prime}$ and $\gamma\colon c\xrightarrow{\sim}c^{\prime}$. We define a map (of sets)
@@
$f(b){f(b^{\prime})}g(c){g(c^{\prime})}$$\scriptstyle{f(\beta)}\scriptstyle{\sim}$$\scriptstyle{g(\gamma)^{-1}\eta^{\prime}f(\beta)}$$\scriptstyle{g(\gamma)}\scriptstyle{\sim}$$\scriptstyle{\eta^{\prime}}$
@@
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(\mathbb{1}^{X},\mathbb{1}^{X})$ $=\prod_{X}\coprod_{X}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(\mathbb{1},\mathbb{1})$
$=\prod_{X}\coprod_{X}\operatorname{Map}_{\operatorname{Fin}}(\ast,\ast)$
@@
Let $G=C_{pq}$ be the cyclic group of order $pq$ for two coprime prime numbers $p$ and $q$. We take the following exhaustive collection of families
$\varnothing\subseteq\{e\}\subseteq\{e,C_{p}\}\subseteq\{e,C_{p},C_{q}\}\subseteq\mathrm{All}.$
@@
(10.5) $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})$$\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Sp}_{G}/\{e\})^{\omega})$$\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Fun}(BG,\operatorname{Sp}^{\omega}))$$\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\mathbb{S}_{p}^{\wedge})^{tG})\times\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\mathbb{S}_{q}^{\wedge})^{tG}).$$\scriptstyle{L^{\omega}_{\{e\}}}$$\scriptstyle{(\widetilde{\Phi}^{e})^{\omega}}$$\scriptstyle{\beta_{0}\circ(\overline{\Phi}^{e})^{\omega}}$$\scriptstyle{\beta_{0}\circ q}$
Using 9.8 and the fact that the corresponding Weyl groups are either $p$-groups or $q$-groups, we see that
$\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Sp}_{G}/\{e\})^{\omega})\simeq(\mathcal{F}\mathrm{in}_{G}/\{e\})^{\operatorname{op}},$
i.e., all separable algebras in $\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Sp}_{G}/\{e\})^{\omega})$ are standard. Using Lemma 7.6, we also find that
@@
We also observe that $(\mathbb{S}_{p}^{\wedge})^{tG}$ and $(\mathbb{S}_{q}^{\wedge})^{tG}$ are indecomposable (argue as in the proof of Theorem 5.5), so the tate categories appearing in the bottom right corner have indecomposable unit. By Proposition 10.2 any object $A\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})$ can be represented by a tuple
@@
    an object $A_{1}\in\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Sp}_{G}/\{e\})^{\omega})$;
@@
    an isomorphism $\eta_{1}\colon A_{1}^{\prime}\simeq A_{2}^{\prime}$ between the images of $A_{1}$ and $A_{2}$ in $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\mathbb{S}_{p}^{\wedge})^{tG})$;
@@
    an isomorphism $\eta_{2}\colon A_{1}^{\prime\prime}\simeq A_{2}^{\prime\prime}$ between the images of $A_{1}$ and $A_{2}$ in $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\mathbb{S}_{q}^{\wedge})^{tG})$.